# Bruno: Backpropagation Running Undersampled for Novel device Optimization

### v1

@@ 3.2.3 Benchmark
Table 3: Parameters used for network simulations with different quantization levels and models. The $alpha_{hid/out}$ parameter is the membrane decay constant for the hidden and output layers. The $\beta_{hid/out}$ parameters are the synapse decay constants for the hidden and output layers. The FeLIF Vthr parameter is the firing threshold of the FeLIF neuron.
@@
Table 4: Test accuracy obtained for different quantization levels and network architectures on the neuromorphic Braille letter reading dataset.
Figure 8: Accuracy Braille dataset with different quantization levels and networks. Each plot has different quantization levels (Floating-point, 8, 4, and 3 bits). In each plot, the three different network architectures are shown, trained with different initialisation seeds using the hyperparameters obtained using HPO
@@ 4 Discussion and conclusion
This work presents a method for training high-temporal neuron models that describe the physical accuracy behaviour of analogue hardware. The simulation of these models can exhibit instability when matching application time steps of $\approx$1 ms. The proposed training method allows the implementation of standard techniques, such as BPTT in the training of the FeLIF model, thereby reducing the training time by an order of magnitude and overcoming problems associated with exploding gradients. The FeLIF neuron model was used to validate the training algorithm. The model was trained with and without BRUNO on the Braille reading dataset, a dataset for benchmarking spatio-temporal patterns. The results show that training with BRUNO improves training time, memory requirements, and accuracy. Achieving a reduction of the 0.28% of the peak memory usage for training long temporal sequences ¿1 s. With an accuracy of $91.34\%\pm 0.59$ on the braille dataset. In contrast, training with vanilla BPTT produces gradient optimisation problems that make it impossible to train the network.
In addition to the validation of the training procedure with the vanilla method, benchmarking was performed with feedforward and recurrent LIF networks. This benchmarking was also performed with synaptic quantization for its possible future implementation on hardware with RRAM synapses. Different quantization levels were performed, for 8, 4, and 3 bits, where Hyperparameters Optimization (HPO) was performed to obtain the optimal set of hyperparameters for each network and quantization level. As a result of this benchmark, we show that a feedforward FeLIF network has comparable results to a feedforward LIF network with $91.34\%\pm 0.59$ and $91.44\%\pm 0.73$ respectively. However, for the quantization network, the FeLIF outperforms the feedforward LIF network and has similar performance with a recurrent LIF network with 3-bit quantization with $74.73\%\pm 1.37$ and $74.23\%\pm 2.48$ respectively. In contrast, the mean accuracy obtained for the feedforward network is $40.48\%\pm 27.84$, which could be due to a high learning rate for the quantized weights, since the network started to lose performance drastically on the last epochs.
BRUNO opens up the possibility of simulating and training neural models with complex dynamics that can be used for spatio-temporal pattern recognition. These neural models have the potential to be integrated into hardware with synaptic quantization, which could provide performance comparable to that of a recurrent LIF network, while requiring fewer resources and trainable parameters.
@@ Acknowledgements
We thank Pedro F. Ortiz Costa and Dr. Francesco Malandrino for their support and fruitful discussions. We thank the Center for Information Technology of the University of Groningen for their support and for providing access to the Hábrók high performance computing cluster. This work was partially supported by the European Research Council (ERC) through the European’s Union Horizon Europe Research and Innovation Programme under Grant Agreement No 101042585. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The authors would like to acknowledge the financial support of the CogniGron research center and the Ubbo Emmius Funds (Univ. of Groningen).

### v2

@@ 3.2.3 Benchmark
Table 4: Parameters used for network simulations with different quantisation levels and models. The $alpha_{hid/out}$ parameter is the membrane decay constant for the hidden and output layers. The $\beta_{hid/out}$ parameters are the synapse decay constants for the hidden and output layers. The FeLIF Vthr parameter is the firing threshold of the FeLIF neuron.
Quant. Model $\alpha_{hid}$ $\beta_{hid}$ $\alpha_{out}$ $\beta_{out}$ LR1 FeLIF Vthr
3 LIF 0.662 0.703 0.565 0.696 $3.429\times 10^{-2}$ -
RLIF 0.603 0.310 0.354 0.295 $1.478\times 10^{-3}$ -
FeLIF 0.468 0.735 - - $7.840\times 10^{-4}$ 3.039
4 LIF 0.226 0.245 0.865 0.834 $3.276\times 10^{-3}$ -
RLIF 0.599 0.710 0.390 0.349 $6.055\times 10^{-3}$ -
FeLIF 0.456 0.322 - - $3.716\times 10^{-3}$ 2.544
8 LIF 0.230 0.302 0.901 0.591 $1.283\times 10^{-2}$ -
RLIF 0.186 0.537 0.347 0.846 $6.766\times 10^{-3}$ -
FeLIF 0.682 0.662 - - $4.387\times 10^{-3}$ 2.928
FP LIF 0.959 0.202 0.716 0.600 $3.006\times 10^{-3}$ -
RLIF 0.343 0.406 0.764 0.874 $6.174\times 10^{-3}$ -
FeLIF 0.299 0.147 - - $2.628\times 10^{-3}$ 3.388
  • 1
    Learning rate
Table 5: Test accuracy obtained for different quantisation levels and network architectures on the neuromorphic Braille letter reading dataset.
@@
Figure 10: Accuracy on the Braille dataset with different quantisation levels and networks. Each plot has different quantisation levels (Floating-point, 8, 4, and 3 bits). In each plot, the three different network architectures are shown, trained with different initialisation seeds using the hyperparameters obtained using HPO
@@ 4 Discussion and conclusion
This work presents a method for training non-linear, high-temporal neuron models derived from the physical mechanisms and behaviour of analogue hardware. Due to the physics-based nature, the simulation of these models can exhibit instability when matching application time steps of $\approx$1 ms. The proposed training method is comparable to standard techniques, such as BPTT, in the training of the FeLIF model, thereby reducing the training time and memory consumption. The FeLIF neuron model was used to validate the training algorithm. The model was trained with and without BRUNO on the JSB dataset for music prediction and the Braille reading dataset, both datasets used for benchmarking spatio-temporal patterns. The results show that training with BRUNO improves training time, memory requirements, and accuracy. Achieving a reduction between 97-99% of the peak memory usage and 50-60% for reduction in training time, compared to BPTT. With an accuracy of $91.34\%\pm 0.59$ on the braille dataset. By contrast, training with BPTT-C involves a trade-off in terms of computational cost, with increased training time and reduced memory usage.
Benchmarking with feedforward and recurrent LIF networks was performed in addition to the validation of the training procedure with BPTT and BPTT-C methods. This benchmark was also performed with synaptic quantisation for its possible future implementation on hardware with RRAM synapses. Different quantisation levels were performed, for 8, 4, and 3 bits, where Hyperparameter Optimization (HPO) was performed to obtain the optimal set of hyperparameters for each network and quantisation level. As a result of this benchmark, we show that a feedforward FeLIF network has comparable results to a feedforward LIF network with $91.34\%\pm 0.59$ and $91.44\%\pm 0.73$ respectively. However, for the quantisation network, the FeLIF outperforms the feedforward LIF network and has similar performance with a recurrent LIF network with 3-bit quantisation with $74.73\%\pm 1.37$ and $74.23\%\pm 2.48$ respectively. In contrast, the mean accuracy obtained for the feedforward network is $40.48\%\pm 27.84$, which could be due to a high learning rate for the quantised weights, since the network started to lose performance drastically on the last epochs.
Thanks to the reduction in resources need for training compared to methods like BPTT or BPTT-C, BRUNO opens up the possibility of simulating and training neural models with complex dynamics that can be used for spatio-temporal pattern recognition. These neural models have the potential to be integrated into hardware with synaptic quantisation, which could provide performance comparable to that of a recurrent LIF network, while not requiring additional recurrent connections, and thus fewer trainable parameters.
@@ Acknowledgements
We thank Pedro F. Ortiz Costa and Dr. Francesco Malandrino, Jun. Prof. David Kappel for their support and fruitful discussions. We thank the Center for Information Technology of the University of Groningen for their support and for providing access to the Hábrók high performance computing cluster. This work was partially supported by the European Union and the European Research Council (ERC) through the European’s Union Horizon Europe Research and Innovation Programme under Grant Agreements No 101042585 and 101207193. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union, the European Research Council or the European Research Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. The authors would like to acknowledge the financial support of the CogniGron research center and the Ubbo Emmius Funds (Univ. of Groningen).